Differentiate from first principle y=tanx


1
Expert's answer
2021-10-15T09:50:49-0400
"(\tan x)'=\lim\limits_{h\to 0}\dfrac{\tan(x+h)-\tan x}{h}"

"=\lim\limits_{h\to 0}\dfrac{\dfrac{\sin(x+h)}{\cos(x+h)}-\dfrac{\sin x}{\cos x}}{h}"

"=\lim\limits_{h\to 0}\dfrac{\dfrac{\sin(x+h)\cos x-\sin x\cos(x+h)}{\cos(x+h)\cos x}}{h}"

"=\lim\limits_{h\to 0}\dfrac{\sin(x+h-x)}{h\cos(x+h)\cos x}"

"=\lim\limits_{h\to 0}\dfrac{\sin(h)}{h}\lim\limits_{h\to 0}\dfrac{1}{\cos(x+h)\cos x}"

"=1(\dfrac{1}{\cos(x+0)\cos x})"

"=\dfrac{1}{\cos^2 x}"

"(\tan x)'=\dfrac{1}{\cos^2 x}"